OpenStudy (twistrose):

Solve for z. 9z+10=1−3(5−3z) z=−4z=−4 z=−2z=−2 all real numbers no solution

OpenStudy (rcp031):

9z+10=1−3(5−3z) 9z+10=1-(15-9z) 9z+10=-14+9z Solution is No solution

OpenStudy (rcp031):

Given equation is invalid.

OpenStudy (twistrose):

OK thank you

OpenStudy (quickstudent):

Yeah I got the same.